5*4a+6(5a+2)
Please help me figure this out.

Answer by elima(1433)

5*4a+6(5a+2)
First do the parenthesis, distribute the 6: 6*5a+6*2 = 30a+12
Now you have:
5*4a+30a+12
next multiply- 5*4a = 20a
now you have:
20a + 30a + 12
combine like terms- 20a + 30a = 50a
now you have:
50a + 12; that is the final answer because you can not add 50 and 12 together; they are not like terms!